# Models That Prove Their Own Correctness

### v2

@@ Abstract
The theoretical framework and results are complemented by experiments on an arithmetic capability: computing the greatest common divisor (GCD) of two integers. Our learning method is used to train a Self-Proving transformer that computes the GCD and proves the correctness of its answer.
@@ 1 Introduction
  • We empirically study TL and Annotated-TL (ATL) for training Self-Proving transformers that compute the Greatest Common Divisor (GCD) of two integers. Table 2 demonstrates the efficacy of our methods, with additional experiments in Section 5. Our results may be of independent interest for research on the arithmetic capabilities of transformers (e.g. Charton 2024; Lee et al. 2024). Code, data and models will be made available upon publication.
@@
Learning method Correctness Verifiability
GPT (baseline) $99.8\%$
GPT+TL $98.8\%$ $60.3\%$
GPT+ATL $98.6\%$ $96.7\%$
Table 2: Self-Proving transformers computing the GCD. We train a 6.7M parameter GPT to compute the GCD of two integers sampled log-uniformly from $[10^{4}]$. Vanilla GPT correctly generates the GCD for almost all inputs, but does not prove correctness to a simple verification algorithm. GPT trained with Transcript Learning (GPT+TL) proves its answer 60.3% of the time; training with Annotated Transcript Learning (GPT+ATL) increases this to 96.7%. See Section 5 for details.
@@ 2 Related Work
This paper is situated at the intersection of machine learning (ML) theory and Interactive Proof systems (IPs). We briefly discuss recent relevant work from these literatures.
@@
IPs have found numerous applications in ML towards a diverse set of goals. Anil et al. (2021) introduce Prover–Verifier Games, a game-theoretic framework for learned provers and verifiers. Wäldchen et al. (2024) cast the problem of model interpretability as a Prover–Verifier interaction between a learned feature selector and a learned feature classifier. Debate systems (Condon et al., 1995), a multiprover variant of IPs, were considered for aligning models with human values (Irving et al., 2018; Brown-Cohen et al., 2023). In such Debate systems, two competing models are each given an alleged answer $y\neq y^{\prime}$, and attempt to prove the correctness of their answer to a (human or learned) judge. Lastly, Murty et al. (2023) define Pseudointelligence: a model learner $L_{M}$ and an evaluator learner $L_{E}$ are each given samples from a ground-truth; $L_{M}$ learns a model of the ground-truth, while $L_{E}$ learns an evaluator of such models; the learned evaluator then attempts to distinguish between the learned model and the ground-truth in a Turing Test-like interaction.
@@
AlphaGeometry (Trinh et al., 2024) is capable of formally proving olympiad-level geometry problems; Gransden et al. (2015); Polu and Sutskever (2020); Yang et al. (2023) and others train models to produce proofs in Coq, Metamath and Lean (de Moura et al., 2015); FunSearch (Romera-Paredes et al., 2024) evolves LLM-generated programs by systematically evaluating their correctness. Indeed, all of these can be cast as Self-Proving models developed for specific proof systems. Meanwhile, this work defines and studies the class of such models in general. Several works (e.g. Welleck et al. 2022) consider models that generate natural language proofs or explanations, which are fundamentally different from formal proofs (or provers) verified by an algorithm.
@@
In Section 5 we train and evaluate Self-Proving transformers to generate the GCD of two integers and prove its correctness to a verifier. These experiments leverage a long line of work on neural models of arithmetic tasks originating with Siu and Roychowdhury (1992). Of particular relevance is the recent paper of Charton (2024), who trains transformers to generate the GCD—without a proof of correctness. We benefit from conclusions suggested in their work and start from a similar (scaled-down) experimental setup. Our main challenge (obtaining Self-Proving models) is overcome by introducing Annotated Transcript Learning (ATL).
@@ 3.1 Preliminaries
Let $\Sigma$ be a set of finite tokens and $\Sigma^{\ast}$ denote the set of finite sequences of such tokens. We consider sequence-to-sequence models $F_{\theta}\colon\Sigma^{\ast}\to\Sigma^{\ast}$, which are total functions that produce an output for each possible sequence. A model is parameterized by a real-valued, finite dimensional vector $\theta$. We consider models as randomized functions, meaning that $F_{\theta}(x)$ is a random variable over $\Sigma^{\ast}$, of which samples are denoted by $y\sim F_{\theta}(x)$.
Before we can define models that prove their own correctness, we must first define correctness. Correctness is defined with respect to an input distribution $\mu$ over $\Sigma^{\ast}$, and a ground-truth $F^{\ast}$ that defines correct answers. For simplicity of presentation, we focus on the case that each input $x\in\Sigma^{\ast}$ has exactly one correct output $F^{\ast}(x)\in\Sigma^{\ast}$, and a zero-one loss function on outputs (the general case is left for future work). The fundamental goal of machine learning can be thought of as learning a model of the ground truth $F^{\ast}$. Formally,
@@
Fix a soundness error $s\in(0,1)$, a finite set of tokens $\Sigma$ and a ground truth $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$. A verifier $V$ (in an Interactive Proof) for $F^{\ast}$ is a probabilistic polynomial-time algorithm that is given explicit inputs $x,y\in\Sigma^{\ast}$ and black-box (oracle) query access to a prover $P$ [^2] We intentionally write $P$ rather than $P_{\theta}$: Interactive Proofs are defined with respect to all possible provers, not just parameterized ones. It interacts with $P$ over $R$ rounds (see Figure 1) and outputs a decision $V,P\right>(x,y)\in\{0,1\}$. Verifier $V$ satisfies the following two guarantees:
@@
As depicted in Figure 1, each of the model’s answers depends on all previous queries and answers in the interaction. This captures the setting stateful models, e.g. a session with a chatbot.
Towards defining Self-Proving models (Section 3.2), let us observe the following. Completeness and soundness are worst-case guarantees, meaning that they hold for all possible inputs $x\in\Sigma^{\ast}$. In particular, completeness implies that for all $x\in\Sigma^{\ast}$, the honest prover $P^{\ast}$ convinces $V$ of the correctness of $F^{\ast}(x)$; in classical proof systems there is no guarantee that an “almost honest” prover can convince the verifier (cf. Paradise (2021)). Yet, if we are to learn a prover $P_{\theta}$, we cannot expect it to agree with $P^{\ast}$ perfectly, nor can we expect it to always output $F^{\ast}(x)$. Indeed, Self-Proving models will have a distributional guarantee with respect to inputs $x\sim\mu$.
@@ 3.2 Self-Proving models
We define the Verifiability of a model $P_{\theta}$ with respect to an input distribution $\mu$ and a verifier $V$. Intuitively, Verifiability captures the ability of the model to prove the correctness of its answer $y\sim P_{\theta}(x)$, when the input $x$ is sampled from $\mu$. We call models capable of proving their own correctness as Self-Proving models.
@@
The benefit of Verifiability over correctness is captured by the following scenario. Alice wishes to use a model $P_{\theta}$ to compute some functionality $F^{\ast}$ on an input $x_{0}$ in a high risk setting. Alice generates $y_{0}\sim P_{\theta}(x_{0})$. Should Alice trust that $y_{0}$ is correct? If Alice has a held-out set of labeled samples, she can estimate $P_{\theta}$’s average correctness on $\mu$. Unfortunately, (average) correctness provides no guarantee regarding for the correctness of the particular $(x_{0},y_{0})$ that Alice has in hand. If, however, Alice has access to a verifier $V$ for which $P_{\theta}$ is Self-Proving, then she can trust the model on an input-by-input (rather than average-case) basis: Alice can execute $V$ on $(x_{0},y_{0})$ and black-box access to $P_{\theta}$. Soundness of $V$ guarantees that if $y_{0}$ is incorrect, then $V$ rejects with high probability, in which case Alice should either generate $P_{\theta}(x_{0})$ again—or find a better model.
@@ 4 Learning Self-Proving Autoregressive Models
Our focus is on autoregressive sequence-to-sequence (Self-Proving) models $P_{\theta}$ (Elman, 1990). Such models generate their output by recursively prompting a randomized sampling from a base distribution $p_{\theta}$ over tokens $\Sigma$. For an input $z\in\Sigma^{\ast}$, the output $w\sim P_{\theta}(z)$ is generated as follows:
@@
  • Let $j=1$. While $w_{j}$ is not the end-of-sequence token $\mathtt{EOS}\in\Sigma^{\ast}$:
@@
For any $z\in\Sigma^{\ast}$, it is useful to consider the vector of log-proabilities over $\Sigma$, denoted by $\log p_{\theta}(z)\in\mathbb{R}^{|\Sigma|}$. We assume that each coordinate in this vector is differentiable with respect to $\theta$.
@@
Section 4.1 introduces Transcript Learning (TL), a learning algorithm that overcomes the exploration problem mentioned in the second point under the assumption that the learner has access to transcripts of interactions in which the verifier accepts. We prove convergence bounds for TL (Section A.1) and analyze it experimentally (Section 5).
@@ 4.1 Transcript Learning
We present an algorithm for learning Self-Proving models which uses access to a distribution of accepting transcripts. This is a reasonable assumption to make when the honest prover $P^{\ast}$ (see Definition 3.2) is efficient, as in the case in Doubly-Efficient Interactive Proof systems as defined by Goldwasser et al. (2015) and developed in other theoretical (e.g. Goldreich and Rothblum 2018) and applied (e.g. Zhang et al. 2021) works. In this case, an honest prover $P^{\ast}$ can be run by the learner during training to collect accepting transcripts without incurring heavy computational cost.
@@
In more detail, Transcript Learning requires access to an (honest) transcript generator $\mathcal{T}^{\ast}$. Given an input $x$, the generator $\mathcal{T}^{\ast}(x)$ samples a sequence $P^{\ast}\!(x)\pi^{\ast}\in\Sigma^{\ast}$ such that $\pi^{\ast}$ is an accepted transcript. The generator is autoregressive, meaning that for any prefix of an accepted transcript $\pi^{\ast}_{\leq t}\in\Sigma^{t}$, the learner has access to the distribution over next tokens $\mathcal{T}^{\ast}(\pi_{\leq t})\in\Sigma$ [^4] Formally, if the generator is prompted with any string that cannot be completed to an accepted transcript, it outputs a dummy symbol $\bot\in\Sigma$.
Transcript Learning (TL) trains a Self-Provable model by autoregressively optimizing towards generating accepting transcripts. It is described in Algorithm 1. At a very high level, it works by repeatedly sampling $x\sim\mu$ and transcript $y^{\ast}\pi^{\ast}\sim\mathcal{T}^{\ast}(x)$, and updating the logits $\log p_{\theta}$ towards agreeing with $y^{\ast}\pi^{\ast}$ via Gradient Ascent. We prove that, under certain conditions, it is expected to output a Self-Provable model.
Theorem 4.1 (Theorem A.5, informal).
@@
is concave and differentiable in $\theta$. For any $\varepsilon>0$, let $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ be upper-bounds such that the following conditions hold.
@@
Denote by $\bar{\theta}$ the output of TL running for number of iterations
@@
The proof (Appendix A) goes by reduction to Stochastic Gradient Descent (SGD). We show that the learner can use only its available tools—sampling honest transcripts, emulating the verifier, and differentiating the logits—to estimate the agreement gradient $\nabla A(\theta)$. Since the agreement $A(\theta)$ lower bounds the Verifiability of $P_{\theta}$, the former can be used as a surrogate objective for the latter.
@@
Quantitatively, the efficiency of TL is captured by the number of iterations $N$. It is desirable to minimize $N$, which is also the number of samples needed from the distribution $\mu$ and the transcript generator $\mathcal{T}^{\ast}$. Like the conditions on the theorem, the bound on $N$ can too be decomposed into two factors: The right factor is the complexity of SGD ($B_{\mathrm{Norm}}^{2}B_{\mathrm{Lip}}^{2}/\varepsilon^{2}$), and the left factor $O(C^{2})$ is the communication complexity of the proof system. Minimizing communication complexity has been an overarching goal in the study of proof systems (e.g. Goldreich and Håstad 1998; Goldreich et al. 2002; Reingold et al. 2021). Theorem 4.1 formally shows the benefit of communication-efficient proof systems in the context of Self-Proving models.

### v3

@@ Abstract
The theoretical framework and results are complemented by experiments on an arithmetic capability: computing the greatest common divisor (GCD) of two integers. Our learning method is used to train a Self-Proving transformer that computes the GCD and proves the correctness of its answer. Our code is available at https://github.com/orrp/self-proving-models.
@@ 1 Introduction
  • We empirically study TL and Annotated-TL (ATL) for training Self-Proving transformers that compute the Greatest Common Divisor (GCD) of two integers. Table 2 demonstrates the efficacy of our methods, with additional experiments in Section 5. Our results may be of independent interest for research on the arithmetic capabilities of transformers (e.g. Charton 2024; Lee et al. 2024). Code, data and models are available at https://github.com/orrp/self-proving-models.
@@
Learning method Correctness (%) Verifiability (%)
GPT (baseline) 99.8 -
GPT+TL 98.8 60.3
GPT+TL+RLVF 98.9 78.3
GPT+Annotated TL 98.6 96.0
Table 2: Self-Proving transformers computing the GCD. We train a 6.3M parameter GPT to compute the GCD of two integers sampled log-uniformly from $[10^{4}]$. Vanilla GPT correctly generates the GCD for almost all inputs, but does not prove correctness to a simple verification algorithm. GPT trained with Transcript Learning (GPT+TL) proves its answer 60.3% of the time; adding Reinforcement Learning from Verifier Feedback (+RLVF) increases this to 78.3%; training with Annotated Transcript Learning (GPT+ATL) gives the highest Verifiability score of 96%. See Section 5 for details.
@@ 2 Related Work
This paper is situated at the intersection of machine learning (ML) and Interactive Proof systems (IPs). We briefly discuss recent relevant work from these literatures.
@@
IPs have found numerous applications in ML towards a diverse set of goals. Anil et al. (2021) introduce Prover–Verifier Games (PVGs), a game-theoretic framework for learned provers and learned verifiers. Since our paper initially appeared, PVGs were further investigated in at least two subsequent works: Hammond and Adam-Day (2024) study multi-prover and Zero Knowledge variants of PVGs. Additionally, Kirchner et al. (2024) successfully utilize PVGs towards obtaining human-legible outputs from LLMs. Notably, they require a relaxed completeness guarantee of their learned proof system—this requirement is the same as our Definition 3.4 of Self-Proving models.
Beyond PVGs, Wäldchen et al. (2024) cast the problem of model interpretability as a Prover–Verifier interaction between a learned feature selector and a learned feature classifier. Debate systems (Condon et al., 1995), a multiprover variant of IPs, were considered for aligning models with human values (Irving et al., 2018; Brown-Cohen et al., 2023). In such Debate systems, two competing models are each given an alleged answer $y\neq y^{\prime}$, and attempt to prove the correctness of their answer to a (human or learned) judge. Lastly, Murty et al. (2023) define Pseudointelligence: a model learner $L_{M}$ and an evaluator learner $L_{E}$ are each given samples from a ground-truth; $L_{M}$ learns a model of the ground-truth, while $L_{E}$ learns an evaluator of such models; the learned evaluator then attempts to distinguish between the learned model and the ground-truth in a Turing Test-like interaction.
@@
AlphaGeometry (Trinh et al., 2024) is capable of formally proving olympiad-level geometry problems; Others have trained models to produce proofs in Gransden et al. (2015); Polu and Sutskever (2020) and others train models to produce proofs in Coq (Gransden et al., 2015), Metamath (Polu and Sutskever, 2020), Lean (Yang et al., 2023), or manually-defined deduction rules (Tafjord et al., 2020); FunSearch (Romera-Paredes et al., 2024) evolves LLM-generated programs by systematically evaluating their correctness. Indeed, all of these can be cast as Self-Proving models developed for specific proof systems. Meanwhile, this work defines and studies the class of such models in general. Several works (e.g. Welleck et al. 2022) consider models that generate natural language proofs or explanations, which are fundamentally different from formal proofs (or provers) verified by an algorithm.
@@
In Section 5 we train and evaluate Self-Proving transformers to generate the GCD of two integers and prove its correctness to a verifier. These experiments leverage a long line of work on neural models for arithmetic tasks originating with Siu and Roychowdhury (1992), and in particular modular arithmetic, which is known to be challenging (Palamas, 2017). Of particular relevance is the recent paper of Charton (2024), who trains transformers to generate the GCD—without a proof of correctness. We benefit from conclusions suggested in their work and start from a similar (scaled-down) experimental setup. Our main challenge (obtaining Self-Proving models) is overcome by introducing Annotated Transcript Learning (ATL).
@@ 3.1 Preliminaries
Let $\Sigma$ be a finite set of tokens and $\Sigma^{\ast}$ denote the set of finite sequences of such tokens. We consider sequence-to-sequence models $F_{\theta}\colon\Sigma^{\ast}\to\Sigma^{\ast}$, which are total functions that produce an output for each possible input sequence. A model is parameterized by a real-valued, finite dimensional vector $\theta$. We consider models as randomized functions, meaning that $F_{\theta}(x)$ is a random variable over $\Sigma^{\ast}$, of which samples are denoted by $y\sim F_{\theta}(x)$.
Before we can define models that prove their own correctness, we must first define correctness. Correctness is defined with respect to an input distribution $\mu$ over $\Sigma^{\ast}$, and a ground-truth $F^{\ast}$ that defines correct answers. For simplicity of presentation, we focus on the case that each input $x\in\Sigma^{\ast}$ has exactly one correct output $F^{\ast}(x)\in\Sigma^{\ast}$, and a zero-one loss function on outputs (the general case is deferred to Appendix A). The fundamental goal of machine learning can be thought of as learning a model of the ground-truth $F^{\ast}$. Formally,
@@
Fix a soundness error $s\in(0,1)$, a finite set of tokens $\Sigma$ and a ground-truth $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$. A verifier $V$ (in an Interactive Proof) for $F^{\ast}$ is a probabilistic polynomial-time algorithm that is given explicit inputs $x,y\in\Sigma^{\ast}$ and black-box (oracle) query access to a prover $P$ [^2] We intentionally write $P$ rather than $P_{\theta}$: Interactive Proofs are defined with respect to all possible provers, not just parameterized ones. It interacts with $P$ over $R$ rounds (see Figure 1) and outputs a decision $V,P\right>(x,y)\in\{\mathrm{reject},\mathrm{accept}\}$. Verifier $V$ satisfies the following two guarantees:
@@
The efficiency of an interactive proof is usually measured with respect to four parameters: the round complexity $R$, the communication complexity (the overall number of bits transferred during the interaction), $P^{\ast}$’s efficiency and $V$’s efficiency. These complexity measures scale with the computational complexity of computing the ground-truth $F^{\ast}$. For example, an interactive proof for a complex $F^{\ast}$ may require multiple rounds of interaction.
@@
As depicted in Figure 1, each of the model’s answers depends on all previous queries and answers in the interaction. This captures the setting of stateful models, e.g. a session with a chatbot.
Towards defining Self-Proving models (Section 3.2), let us observe the following. Completeness and soundness are worst-case guarantees, meaning that they hold for all possible inputs $x\in\Sigma^{\ast}$. In particular, completeness implies that for all $x\in\Sigma^{\ast}$, the honest prover $P^{\ast}$ convinces $V$ of the correctness of $F^{\ast}(x)$; in classical proof systems there is no guarantee that an “almost honest” prover can convince the verifier (cf. Paradise 2021). Yet, if we are to learn a prover $P_{\theta}$, we cannot expect it to agree with $P^{\ast}$ perfectly, nor can we expect it to always output $F^{\ast}(x)$. Indeed, Self-Proving models will have a distributional guarantee with respect to inputs $x\sim\mu$.
@@ 3.2 Self-Proving models
We define the Verifiability of a model $P_{\theta}$ with respect to an input distribution $\mu$ and a verifier $V$. Intuitively, Verifiability captures the ability of the model to prove the correctness of its answer $y\sim P_{\theta}(x)$, when the input $x$ is sampled from $\mu$. We refer to models capable of proving their own correctness as Self-Proving models. Notice that, as in Definition 3.2, the verifier is fixed and agnostic to the choice of the Self-Proving model.
@@
The benefit of Verifiability over correctness is captured by the following scenario. Alice wishes to use a model $P_{\theta}$ to compute some functionality $F^{\ast}$ on an input $x_{0}$ in a high risk setting. Alice generates $y_{0}\sim P_{\theta}(x_{0})$. Should Alice trust that $y_{0}$ is correct? If Alice has a held-out set of labeled samples, she can estimate $P_{\theta}$’s average correctness on $\mu$. Unfortunately, (average) correctness provides no guarantee regarding the correctness of the particular $(x_{0},y_{0})$ that Alice has in hand. If, however, Alice has access to a verifier $V$ for which $P_{\theta}$ is Self-Proving, then she can trust the model on an input-by-input (rather than average-case) basis: Alice can execute $V$ on $(x_{0},y_{0})$ and black-box access to $P_{\theta}$. Soundness of $V$ guarantees that if $y_{0}$ is incorrect, then $V$ rejects with high probability, in which case Alice should either generate $P_{\theta}(x_{0})$ again—or find a better model.
@@ 4 Learning Self-Proving Autoregressive Models
Our focus is on autoregressive sequence-to-sequence (Self-Proving) models $P_{\theta}$. Such models generate their output by recursively prompting a randomized sampling from a base distribution $p_{\theta}$ over tokens $\Sigma$. For an input $z\in\Sigma^{\ast}$, the output $w\sim P_{\theta}(z)$ is generated as follows:
@@
  • Let $j=1$. While $w_{j}$ is not the end-of-sequence token $\mathtt{EOS}\in\Sigma$:
@@
For any $z\in\Sigma^{\ast}$, it is useful to consider the vector of log-probabilities over $\Sigma$, denoted by $\log p_{\theta}(z)\in\mathbb{R}^{|\Sigma|}$. We assume that each coordinate in this vector is differentiable with respect to $\theta$.
@@
Section 4.1 introduces Transcript Learning (TL), a learning algorithm that overcomes the exploration problem mentioned in the second point under the assumption that the learner has access to transcripts of interactions in which the verifier accepts. We prove convergence bounds for TL (Section B.1) and analyze it experimentally (Section 5).
@@ 4.1 Transcript Learning
We present an algorithm for learning Self-Proving models which uses access to a distribution of accepting transcripts. This is a reasonable assumption to make when the honest prover $P^{\ast}$ (see Definition 3.2) is efficient, as in the case of public-coin Doubly-Efficient Interactive Proof systems as defined by Goldwasser et al. (2015) and developed in other theoretical (e.g. Goldreich and Rothblum 2018) and applied (e.g. Zhang et al. 2021) works. In this case, an honest prover $P^{\ast}$ can be run by the learner during training to collect accepting transcripts without incurring heavy computational cost. Alternatively, the learner may collect a dataset of accepting transcripts prior to learning (see Figure 4 in Appendix B).
@@
In more detail, Transcript Learning (TL, Algorithm 1) requires access to an (honest) transcript generator $\mathcal{T}^{\ast}$. Given an input $x$, the generator $\mathcal{T}^{\ast}(x)$ samples a sequence $P^{\ast}\!(x)\pi^{\ast}\in\Sigma^{\ast}$ such that $\pi^{\ast}$ is an accepted transcript. TL trains a Self-Provable model by autoregressively optimizing towards generating accepting transcripts. At a very high level, it works by repeatedly sampling $x\sim\mu$ and transcript $y^{\ast}\pi^{\ast}\sim\mathcal{T}^{\ast}(x)$, and updating the logits $\log p_{\theta}$ towards agreeing with $y^{\ast}\pi^{\ast}$ via Gradient Ascent. We prove that, under certain conditions, it is expected to output a Self-Provable model.
Theorem 4.1 (Theorem B.5, informal).
@@
is concave in $\theta$. For any $\varepsilon>0$, let $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ be upper-bounds such that the following conditions hold.
@@
Denote by $\bar{\theta}$ the output of Transcript Learning (Algorithm 1) running for $N$ interations, where
@@
The proof (Appendix B) goes by reduction to Stochastic Gradient Descent (SGD). We show (Lemma B.4) that the learner can use its only available tools—sampling honest transcripts, emulating the verifier, and differentiating the logits—to optimize the agreement $A(\theta)$. Specifically, this is done by accumulating gradients from the cross-entropy loss computed at each token. Since $A(\theta)$ lower bounds the Verifiability of $P_{\theta}$, the former can be used as a surrogate for the latter.
@@
Quantitatively, the efficiency of TL is captured by the number of iterations $N$. It is desirable to minimize $N$, which is also the number of samples needed from the distribution $\mu$ and the transcript generator $\mathcal{T}^{\ast}$. The bound on $N$ in Equation 2 can be decomposed into the complexity of SGD ($B_{\mathrm{Norm}}^{2}B_{\mathrm{Lip}}^{2}/\varepsilon^{2}$), and communication complexity of the proof system $O(C^{2})$. Minimizing communication complexity has been an overarching goal in the study of proof systems (e.g. Goldreich and Håstad 1998; Goldreich et al. 2002; Reingold et al. 2021). Theorem 4.1 formally shows the benefit of communication-efficient proof systems in the context of Self-Proving models.